Hello, Im having problems with setting this up correctly.
Hassan is designing a garden in the shape of a right triangle. One leg of the triangle is 8 ft long. The other leg and the hypotenuse are consecutive odd integers. How long are the other two sides of the garden.
so I have
hypotenuse =x+2
leg 1 = x
leg 2 = 8
is x^2+8^2=(x+2)^2 on the right track? im not sure if im on the right course or where to take it from here
any help is appreciated, thank you in advance

Found 3 solutions by mananth, richwmiller, stanbon:
Answer by mananth(16946)

hypotenuse =x+2
leg 1 = x
leg 2 = 8
is x^2+8^2=(x+2)^2

on the right track
x^2+64=x^2+4x+4
64=4x+4
60=4x
x=15
other leg
hypotenuse is 17
8,15.17 triplet

Answer by richwmiller(17219)

x^2+8^2=(x+2)^2
looks good
x=15 x+2=17

Answer by stanbon(75887)

One leg of the triangle is 8 ft long. The other leg and the hypotenuse are consecutive odd integers. How
---------
leg:: 2x-1
hypo: 2x+1
----------
leg^2 + leg^2 = hypotenuse^2
------
8^2 + (2x-1)^2 = (2x+1)^2
-----------------------------------
64 + 4x^2-4x+1 = 4x^2+4x+1
-----------------------------------
64 = 8x
x = 8
-----
leg:: 2x-1 = 15
Hypo:: 2x+1 = 17
-----
Checking:
8^2 + 15^2 = 17^2
64 + 225 = 289
289 = 289
------------------
Cheers,
Stan H.
==================